# Blinded by Generated Contexts: How Language Models Merge Generated and Retrieved Contexts for Open-Domain QA?

### v4

@@ 4.3 Consistent Preferences Across Retrievers
The experiments conducted above have established the presence of the bias in LLMs against contexts retrieved by Contriever, a dense retrieval model. This section further explores whether such bias remains consistent across different retrieval models. To this end, we incorporate BM25 as a representative of sparse retrieval techniques, along with human-labeled golden passages, to mimic an ideal retrieval model (Gold). For each retrieval model, we respectively construct AIR context-conflicting datasets based on the NQ development sets from Karpukhin et al. (2020), which include the golden passages. As illustrated in Table 3, the results indicate that LLMs consistently display a pronounced bias in favor of generated contexts regardless of the retrieval model used.
@@ 5 Why LLMs Prefer Generated Contexts
In this section, we investigate why LLMs prefer generated contexts rather than retrieved contexts, from several perspectives: confirmation bias in §5.1, context similarity to the question in §5.2, and context completeness in §5.3.
5.1 Effect of Confirmation Bias
Recently, Xie et al. (2023) identified a confirmation bias in LLMs, wherein they exhibit a preference for contexts consistent with their parametric knowledge (also known as parametric memory) when faced with two conflicting generated contexts. In our experiments discussed in Section 4.1, which involve a single LLM serving both as reader and generator, it is natural to assume that the generated contexts align with the LLM’s parametric knowledge. This assumption leads to a key question: Does the confirmation bias lead to the observed preference for generated contexts in this work?
To investigate the effect of confirmation bias, we designed controlled experiments that disrupt the consistency between generated contexts and LLMs’ parametric knowledge. Overall, we enforce LLMs to make up a counter-memory context $d^{\varrho^{\prime}}_{1}$, which supports a same-type yet different answer compared to the original generated context. Then, we replace the generated context $d^{\varrho}_{1}$ with the counter-memory context $d^{\varrho^{\prime}}_{1}$ and recalculate $\mathrm{DiffGR}$ to assess shifts in preference after excluding confirmation bias, as detailed in Section A.6. Table 4 reveals that LLMs maintain a significant bias to generated contexts when they are inconsistent with LLMs’ parametric knowledge, indicating that confirmation bias does not play a major role in the observed bias. This finding does not negate the presence of confirmation bias, since other factors distinguishing generated from retrieved contexts might exert a more significant influence, potentially overshadowing the effects of confirmation bias.
Notably, GPT-3.5 even shows a stronger bias towards counter-memory contexts on TQA-AIR, a phenomenon also observed in Xie et al. (2023). This is likely attributed to the higher similarity of counter-memory contexts generated by GPT-3.5, on both our dataset and the ConflictQA datasets of Xie et al. (2023), as discussed in Section A.6.2.
Context pair NQ-AIR TQA-AIR GPT-3.5 Llama2-13b GPT-3.5 Llama2-13b Gen vs. Ret 0.7561 0.6747 0.7058 0.6575 Ctr   vs. Ret 0.7342 0.6468 0.8010 0.6596
Table 4: $\mathrm{DiffGR}$ of different input context pairs. “Gen”, “Ret” and “Ctr” respectively represent generated contexts, retrieved contexts, and counter-memory contexts.
@@ 5.2 Effect of Text Similarity
Figure 7: Context-question similarity distribution of generated and retrieved contexts on the union of AIR subsets for different LLMs. Distribution on AIG subsets is presented in Appendix A.7, with similar results.
Figure 8: The $\mathrm{DiffGR}$ in slices with different average $\Delta\operatorname{sim}$. Results for other LLMs are in Appendix A.7.
To further clarify the influence of the observed similarity discrepancies, we rank the samples according to the similarity gap $\Delta\operatorname{sim}$ between generated and retrieved contexts. Then, we divide the dataset into $n$ ($n=5$ here [^2] Similar results and observations are found with other $n$.) slices with an equal number of samples.
@@
Here, $\operatorname{sim}(q,d^{\varrho})$ is the similarity between generated context and question, and $\operatorname{sim}(q,d^{\gamma})$ is for retrieved context. Figure 8 illustrates the relationship between the average $\Delta\operatorname{sim}$ within each slice and the corresponding $\mathrm{DiffGR}$. From the results, we observe a general trend that LLMs exhibit an increased bias to generated contexts on slices with a larger average similarity gap, which indicates that text similarity is a significant factor in the preference for generated contexts. These findings suggest that generated contexts should be applied with greater caution to mitigate the influence of highly relevant but misleading information.
To facilitate understanding why the similarity affects LLMs’ preference, we include some examples in Appendix A.7.4. From these cases, we observe that contexts with higher similarity often support candidate answers more straightforwardly, for instance, by mirroring the phrasing used in the questions. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences.
@@ 5.3 Effect of Context Completeness
Context Length Completeness Similarity Semantic Sentence Jaccard BERTScore Retrieved 107.4 0.114 0.801 Nature 109.7 0.202 0.879 Trunc. 107.4 0.196 0.877 S-Trunc. 105.9 0.193 0.876
Table 5: Average length and similarity of contexts with different completeness (details in Appendix A.2.2).
Context Pair NQ-AIR TQA-AIR GPT-3.5 Llama2-13b GPT-3.5 Llama2-13b Nature vs. Ret 0.6562 0.5987 0.6101 0.5799 Trunc. vs. Ret 0.2536 0.0958 0.3581 0.2042 S-Trunc. vs. Ret 0.2394 0.1071 0.3302 0.1958
@@
To investigate the potential effects of completeness on the observed bias, we conduct controlled experiments that vary the semantic and sentence completeness of generated contexts [^3] We also attempted to control the completeness of retrieved contexts, but it was challenging to isolate it from confounding factors like length. This aspect is left for future work.using the following methods: (a) Truncation (Trunc.) eliminates the length constraints from the generation prompt of Section 3, allowing LLMs to generate extended contexts. These generated contexts are then truncated to match the length of retrieved contexts, thereby simulating both semantic and sentence incompleteness of retrieved contexts. (b) Sentence Truncation (S-Trunc.): Based on the method (a), we truncate generated contexts only at the end of a sentence to preserve the sentence completeness, while simulating the semantic incompleteness.
@@
Table 5 demonstrates that three types of generated contexts have similar average lengths and similarities. This suggests that the influences of similarity and length are mitigated, thereby highlighting the principal disparities in semantic content and sentence completeness.
@@
We evaluate LLMs’ preference between generated versus retrieval context, varying the completeness of generated context, following the same pipeline in Section 4.1. Table 6 presents the $\mathrm{DiffGR}$ with different semantic and sentence completeness in generated contexts. A comparison between “Trunc.” and “S-Trunc.” reveals that sentence completeness does not significantly affect LLMs’ preference for generated contexts. In contrast, comparing “Nature” and “S-Trunc.”, we find a significant increase in bias towards generated contexts, when they are semantically more complete. These findings indicate that LLMs tend to favor contexts with enhanced semantic completeness, underscoring the necessity to investigate improved passage segmentation methods that maintain semantic completeness for current retrieval-augmented LMs.
@@ 6 Conclusion and Future work
In this study, we propose a framework to investigate the underlying mechanisms by which LLMs merge retrieved and generated contexts. Our results reveal a pronounced bias towards generated contexts in several LLMs (GPT 3.5/4 and Llama2-7b/13b). We further identify two key factors that may contribute to this bias: higher similarity between generated contexts and questions, and the semantic incompleteness of retrieved contexts.
Our insights highlight the critical need for advanced integration methods that can validate and leverage information from both sources, moving beyond the current overreliance on generated contexts. Additionally, we find that LLMs display significant sensitivity to the semantic completeness of input contexts. This sensitivity necessitates improved passage segmentation strategies in current retrieval-augmented systems, thereby ensuring the preservation of intended meaning and the maximization of utility. Finally, addressing the challenges posed by highly relevant yet incorrect information generated by LLMs is an important direction for future research. It is crucial to develop methods for detecting and discounting misleading information produced by LLMs, especially as the volume of such content continues to escalate.
Limitation
@@
Appendix A Appendix
A.1 Related work
A.1.1 Generation-Augmented Approaches
Generation-augmented approaches prompt LLMs to generate intermediate contexts for the final response, thereby leveraging their extensive parametric knowledge acquired during the pre-training phase on vast text corpora (Roberts et al., 2020; Petroni et al., 2019). Recent studies have demonstrated the effectiveness of these methods across a range of tasks. Liu et al. (2022) represents an early exploration into augmenting LLMs in commonsense reasoning tasks using knowledge generated by LLMs themselves. Sun et al. (2023); Yu et al. (2022) employ LLMs to produce background documents (or recitations) and subsequently use these documents to enhance LLM performance in knowledge-intensive tasks. Another line of research, known as the chain-of-thought (Wei et al., 2022; Kojima et al., 2022), prompts LLMs to generate intermediate reasoning steps to enhance LLMs’ reasoning abilities. Despite the effectiveness of these methods, the LLM-generated knowledge may contain hallucinations (Chen et al., 2023; Ji et al., 2023) due to LLMs’ outdated memory (De Cao et al., 2021) and limited memory for long-tail knowledge (Kandpal et al., 2023). The LLM-generated inaccuracy information could potentially mislead current retrieval model (Dai et al., 2023) and open-domain question answering systems (Pan et al., 2023).
A.1.2 Retrieval-Augmented Approaches
The retrieval-augmented approaches (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023; Gao et al., 2023) enhance LLMs by incorporating relevant documents from the external corpus. These approaches represent a promising direction for addressing the knowledge limitations of LLMs, such as the need for knowledge updating (Jang et al., 2022) and long-tail knowledge (Kandpal et al., 2023). Early retrieval-augmented methods (Guu et al., 2020; Lewis et al., 2020; Izacard et al., 2022) focused on the joint training of LLMs and retrieval modules to improve their cooperation. With the evolution of general-purpose LLMs, recent studies (Ram et al., 2023; Shi et al., 2023) investigate the strategy of appending relevant documents directly to the input while keeping the LLMs frozen. Despite the effectiveness of these methods, the retrieval-augmented approaches still face challenges due to irrelevant retrieval results and incomplete knowledge coverage Yu et al. (2023); Mallen et al. (2023). These noisy retrieval results can misguide the outcomes of LLMs (Mallen et al., 2023; Yoran et al., 2023; Ren et al., 2023).
A.1.3 Hybrid Approaches and Knowledge Conflicts
Recent works investigate merging retrieved and generated contexts to leverage both parametric knowledge and external knowledge Abdallah and Jatowt (2023); Yu et al. (2022). These combination methods have shown improved performance over those relying solely on a single information source in a fully-supervised setting. Furthermore, (Zhang et al., 2023) proposes an improved method to effectively leverage the two sources of information, especially when conflicts arise. While these works have focused on improving the efficacy of hybrid approaches, the underlying mechanisms by which LLMs process conflicting information from different types of contexts remain underexplored.
Current research on knowledge conflicts in LLMs primarily focuses on two aspects: conflicts between input contexts and LLMs’ parametric memory, and conflicts among the input contexts themselves. Regarding the former, Xie et al. (2023); Chen et al. (2022) find that LLMs are highly receptive to the input contexts rather than their internal memory. Concerning conflicts within multiple input contexts, Chen et al. (2022) demonstrates that LLMs tend to rely on a few most relevant retrieved contexts. Additionally, Xie et al. (2023) reveals a confirmation bias, i.e., LLMs demonstrate a tendency to favor contexts that align with their parametric knowledge when confronted with both supporting and opposing contexts. However, these studies are limited to analyzing context conflicts within a single type of input context. Our work complements these studies by considering conflicts between generated and retrieved contexts and reveals several key factors, such as semantic completeness and text similarity.
A.2 Length Control for Generated Contexts
A.2.1 Length Distribution across LLMs

### v6

@@ 4.3 Consistent Preferences Across Retrievers
The above experiments have demonstrated the bias in LLMs against contexts retrieved by Contriever, a dense retrieval model. This section further explores whether such bias remains consistent across different retrieval models. To this end, we incorporate BM25 as a representative of sparse retrieval techniques, along with human-labeled golden passages, to mimic an ideal retrieval model (Gold). For each retrieval model, we respectively construct AIR context-conflicting datasets based on the NQ dev sets from Karpukhin et al. (2020), which include the golden passages. As illustrated in Table 3, the results indicate that LLMs consistently display a pronounced bias in favor of generated contexts regardless of the retrieval model used.
@@ 5 Why LLMs Prefer Generated Contexts
In this section, we investigate the causes of the observed bias from several perspectives: the effect of parametric knowledge in §5.1, context similarity to the question in §5.2, and context completeness in §5.3. This section primarily presents the result for Llama2-13b and GPT-3.5, while the results for the other LLMs are provided in Appendix B.
Reader Generator NQ (12367) TQA (20150) NQ-AIG NQ-AIR TQA-AIG TQA-AIR GPT 3.5 GPT 3.5 500 457 524 322 GPT 3.5 Llama2-13b 271 665 359 574 Llama2-13b GPT 3.5 1318 553 1889 467 Llama2-13b Llama2-13b 633 841 928 1020
Table 4: The number of data in selected subsets where $a_{\phi}^{\text{llm}}\neq a_{\phi}^{\gamma}\neq a_{\phi}^{\varrho}$. More results are shown in Table 10.
5.1 Effect of Parametric Knowledge
Recently, Xie et al. (2023) demonstrated that LLMs exhibit a bias towards contexts consistent with their parametric knowledge (or memory), a phenomenon termed confirmation bias. This section explores whether the observed bias arises from the potential consistency between generated contexts and parametric knowledge. To investigate this, we select subsets where the answers from “retrieved context”, “generated context”, and “parametric knowledge” were all different from one another.
To achieve the subset, we first establish the LLM’s parametric knowledge about a question using a closed-book QA task, following Xie et al. (2023), which can be expressed as $a_{\phi}^{\text{llm}}=\phi(q)$. Then, we select the cases that satisfy $a_{\phi}^{\text{llm}}\neq a_{\phi}^{\gamma}\neq a_{\phi}^{\varrho}$ from the AIR datasets constructed in §3. Table 4 shows the number of samples after filtering.
The filtered datasets not only exclude the effect of confirmation bias but also help us identify whether LLMs use parametric knowledge or contexts to answer the question. Table 5 illustrates the proportion of LLMs choosing the answer provided by generated contexts, retrieved contexts, or parametric knowledge. We observe two key insights: (i) The proportion of choosing LLMs’ parametric knowledge is very small. This result is consistent with the findings of several previous works (Xie et al., 2023; Chen et al., 2022), which found that LLMs mostly rely on the context even when it conflicts with the parametric knowledge. (ii) LLMs still show a significant preference for generated contexts when excluding the influence of parametric knowledge. This suggests that the confirmation bias is not the primary cause of the observed bias.
To ensure rigorous analysis, we use the filtered data in Table 4 for subsequent experiments to eliminate any potential distractions.
Reader Generator $\rho_{\text{gen}}$ $\rho_{\text{ret}}$ $\rho_{\text{llm}}$ $\mathrm{DiffGR}$ $\mathrm{DiffGR}$ (ori) NQ-AIR GPT 3.5 GPT 3.5 67.83 12.91 0.88 0.68 0.66 GPT 3.5 Llama2-13b 67.22 15.34 1.20 0.63 0.58 Llama2-13b GPT 3.5 62.39 26.76 1.08 0.40 0.42 Llama2-13b Llama2-13b 69.92 18.67 1.43 0.58 0.60 TQA-AIR GPT 3.5 GPT 3.5 72.05 16.15 1.55 0.63 0.61 GPT 3.5 Llama2-13b 70.21 16.38 2.61 0.62 0.62 Llama2-13b GPT 3.5 61.88 26.34 1.71 0.40 0.35 Llama2-13b Llama2-13b 72.55 17.75 1.96 0.61 0.58
Table 5: The columns $\rho_{\text{gen}}$, $\rho_{\text{ret}}$, and $\rho_{\text{llm}}$ represent the proportion (%) of responses matching answers from generated contexts, retrieved contexts, and parametric knowledge, respectively. These values are computed on the filtered subset where $a_{\phi}^{\text{llm}}\neq a_{\phi}^{\gamma}\neq a_{\phi}^{\varrho}$, except $\mathrm{DiffGR}$ (ori), which is computed on the original dataset. Results for other LLMs are shown in Table 11.
@@ 5.2 Effect of Text Similarity
Figure 7: Context-question similarity distribution of generated and retrieved contexts on the union of AIR subsets for different LLMs. More results are shown in Section B.5.
Figure 8: The $\mathrm{DiffGR}$ in slices with different average $\Delta\operatorname{sim}$. Results for other LLMs are in Figure 15.
To further clarify the influence of the observed similarity discrepancies, we rank the samples according to the similarity gap $\Delta\operatorname{sim}$ between generated and retrieved contexts.
@@
Here, $\operatorname{sim}(q,d^{\varrho})$ is the similarity between generated context and question, and $\operatorname{sim}(q,d^{\gamma})$ is for retrieved context. Then, we divide the dataset into $n$ ($n=5$ here [^2] Similar results and observations are found with other $n$.) slices with an equal number of samples. Ensuring that each slice contains an equal number of samples helps to avoid fluctuations caused by sample size variations.
Figure 8 illustrates the relationship between the average $\Delta\operatorname{sim}$ within each slice and the corresponding $\mathrm{DiffGR}$ [^3] We also tried to manipulate the context similarity to investigate its impact, but we found it challenging to instruct LLMs to generate contexts with low similarity to the question while still providing an answer to it.. From the results, we observe a general trend that LLMs exhibit an increased bias to generated contexts on slices with a larger average similarity gap, which indicates that text similarity is a significant factor in the preference for generated contexts. These findings suggest that generated contexts should be applied with greater caution to mitigate the influence of highly relevant but misleading information.
To facilitate understanding why the similarity affects LLMs’ preference, we include some examples in Table 16. From these cases, we observe that contexts with higher similarity often support candidate answers more straightforwardly, for instance, by mirroring the phrasing used in the questions. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences.
@@ 5.3 Effect of Context Completeness
Context Completeness Similarity Length Sentence Semantic Jaccard BERTScore Retrieved 0.1162 0.8552 107.1 Nature 0.1748 0.8730 105.7 S-Trunc. 0.1733 0.8730 105.8 Trunc. 0.1769 0.8736 107.1
Table 6: Average length and similarity of contexts with different completeness on NQ-AIR dataset for GPT-3.5 (more details in Table 14).
Context Pair NQ-AIR TQA-AIR GPT-3.5 Llama2-13b GPT-3.5 Llama2-13b Nature vs. Ret. 0.7519 0.6082 0.6353 0.6207 S-Trunc. vs. Ret. 0.4864 0.2551 0.5802 0.2779 Trunc. vs. Ret. 0.4792 0.2198 0.5663 0.2787
@@
To investigate the potential effects of completeness on the observed bias, we conduct controlled experiments that vary the semantic and sentence completeness of generated contexts [^4] We also tried to vary the completeness of retrieved contexts but found it challenging to isolate it from confounding factors like length. This aspect is left for future work. using the following methods: (a) Truncation (Trunc.) eliminates the length constraints from the generation prompt of §3, allowing LLMs to generate extended contexts. These generated contexts are then truncated to match the length of retrieved contexts, thereby simulating both semantic and sentence incompleteness of retrieved contexts. (b) Sentence Truncation (S-Trunc.): Based on the method (a), we truncate generated contexts only at the end of a sentence to preserve the sentence completeness, while simulating the semantic incompleteness.
@@
To eliminate the interference of similarity factors, we select questions whose three types of generated contexts have nearly equivalent similarity, with BERTScore differences less than 0.05555We also tried several other thresholds for similarity differences, e.g. 0.005, all yielding same conclusions.. Table 6 demonstrates that three types of generated contexts also have similar average lengths after filtering. This means that both the influences of similarity and length are mitigated, thereby highlighting the principal disparities in semantic and sentence completeness.
@@
We evaluate LLMs’ preference between generated versus retrieved context, varying the completeness of generated context, following the same pipeline in §4.1. Table 7 presents the $\mathrm{DiffGR}$ with different semantic and sentence completeness in generated contexts. A comparison between “Trunc.” and “S-Trunc.” reveals that sentence completeness has a very slight impact on LLMs’ preference for generated contexts. In contrast, comparing “Nature” and “S-Trunc.”, we find a significant increase in bias towards generated contexts that are semantically more complete. These findings indicate that LLMs tend to favor contexts with enhanced semantic completeness, underscoring the necessity to investigate improved passage segmentation methods that maintain semantic completeness for current retrieval-augmented LMs.
6 Related Work
6.1 Generation-Augmented Approaches
Generation-augmented methods prompt LLMs to generate intermediate contexts for the final response, thereby leveraging their extensive parametric knowledge acquired during the pre-training phase on vast text corpora (Roberts et al., 2020; Petroni et al., 2019). These generated contexts may encompass various types of knowledge, such as background knowledge (Sun et al., 2023; Yu et al., 2022), commonsense knowledge (Liu et al., 2022), domain-specific knowledge (Feng et al., 2023; Luo et al., 2023), and chain-of-thought reasoning processes (Wei et al., 2022; Kojima et al., 2022). Despite their effectiveness, the LLM-generated knowledge may contain hallucinations (Chen et al., 2023; Ji et al., 2023) due to LLMs’ outdated memory (De Cao et al., 2021) and limited memory for long-tail knowledge (Kandpal et al., 2023). This inaccuracy information could potentially mislead current retrieval model (Dai et al., 2023) and open-domain question answering systems (Pan et al., 2023).
6.2 Retrieval-Augmented Approaches
Retrieval-augmented methods (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023; Gao et al., 2023) enhance LLMs by integrating relevant documents from external corpora, addressing limitations like the need for knowledge updates (Jang et al., 2022) and long-tail knowledge (Kandpal et al., 2023). Early methods (Guu et al., 2020; Lewis et al., 2020; Izacard et al., 2022) focused on joint training of LLMs and retrievers. Recent studies (Ram et al., 2023; Shi et al., 2023) append relevant documents directly to the input while keeping LLMs static. Despite their effectiveness, these methods face challenges with irrelevant retrievals and incomplete knowledge coverage (Yu et al., 2023; Mallen et al., 2023). These noisy retrieval results can misguide LLMs (Mallen et al., 2023; Yoran et al., 2023; Ren et al., 2023).
6.3 Hybrid Approaches and Knowledge Conflicts
Recent works investigate merging retrieved and generated contexts to leverage both sources of knowledge (Abdallah and Jatowt, 2023; Yu et al., 2022; Zhang et al., 2023) and achieve improved performance over those relying solely on a single information source. However, conflicts between diverse sources can impede the effectiveness of information integration Zhang et al. (2023) Current research on knowledge conflicts primarily focuses on two aspects: context-memory conflict and inter-context conflict (Xu et al., 2024). Regarding context-memory conflict, Xie et al. (2023); Chen et al. (2022) find that LLMs are highly receptive to the input contexts rather than their internal memory. Concerning inter-context conflict, Chen et al. (2022) demonstrates that LLMs tend to rely on a few most relevant retrieved contexts. Additionally, Xie et al. (2023) reveals that LLMs favor contexts consistent with their parametric knowledge when confronted with both supporting and opposing contexts. In contrast to these studies which are limited to conflicts within a single type of context, our work further considers conflicts between generated and retrieved contexts and reveals a bias in LLMs.
@@ 7 Conclusion and Future Work
In this study, we propose a framework to investigate the underlying mechanisms by which LLMs merge retrieved and generated contexts. Our results reveal a pronounced bias towards generated contexts in several LLMs (GPT 3.5/4 and Llama2-7b/13b). We further identify two key factors that may contribute to this bias: higher similarity between generated contexts and questions, and the semantic incompleteness of retrieved contexts. Our insights highlight the critical need for advanced integration methods that can validate and leverage information from both sources, moving beyond the current overreliance on generated contexts.
While this study primarily focuses on integrating generated and retrieved contexts, the observed bias also highlights a critical risk of retrieval-augmented LLMs. With LLM-generated content increasingly prevalent on the internet, retrieval results may include generated contexts (Chen and Shu, 2023). As this scenario becomes more common, the observed bias of LLMs toward generated contexts implies their susceptibility to misinformation and malicious attacks generated by LLMs. This raises concerns about the security of retrieval-augmented systems, a critical problem that is gaining attention in recent works (Pan et al., 2023; Zou et al., 2024). Addressing the challenges posed by the widespread presence of generated content, such as the detection of such content, represents a promising direction for future research.
Limitations
@@
Acknowledgements
This work was supported by the National Key R&D Program of China (2022YFB3103700, 2022YFB3103704), the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680202), the Innovation Funding of ICT, CAS (E361120), the National Natural Science Foundation of China (No.62172393), and Major Public Welfare Project of Henan Province (No.201300311200).
Appendix A Detailed Datasets Statistics
A.1 Length Control for Generated Contexts